Pheromones are compounds secreted by the females of many insect species to attract males. One of these compounds contains 80.78 percent C, 13.56 percent H, and 5.66 percent O. A solution of 1.00 g of this pheromone in 8.50 g of benzene freezes at 3.37
Answer #1
?Tf = Kf • mB

Kf for benzene = 5.12

(5.5 - 3.37) = 5.12 x molality

molality = 2.13 / 5.12 = 0.416 moles / Kg

we have 1 g / 8.5 g of solution or 1 X 1000 / 8.5 = 117.6 g / kg of solution

our molality is .416 moles / kg;

so .416 moles = 117.6 g

1 mole = 117.6 / .416 = 282.7g .. molecular weight

mass of C in the compound 80.78 % of 282.7 = 228.4 g or 228 .4 / 12 = 19 moles of C

mass of H = 13.56 % of 282.7 = 38.2 g or 38 moles of H

mass of O = 5.66 % of 282.7 = 16 g or 1 mole

Molecular formula is C19H38O